OpenStudy (anonymous):

Tammy solved these problems. 18 ÷ 6 = 18 ÷ 3 = 18 ÷ 1 = Which observation could Tammy make about the quotients after she had her solutions? A. The quotient decreased as the divisor decreased. B. The quotient increased as the divisor decreased.

OpenStudy (arihangdu):

18/3=6, 18/6=3, and 18/1 = 18

OpenStudy (arihangdu):

Thanks lol💕

OpenStudy (dallascowboys88):

the answer is 6,3,and18